OpenStudy (anonymous):

HELP.....write the equation of a line with the slop, m=3/5, which contains the point (2,3) FILL IN THE BLANKS Y-_____=______(X-______) Please explain

OpenStudy (saifoo.khan):

y - 3 = 3/5 x - 2

OpenStudy (karatechopper):

aww darn it saif i wanted this prob

OpenStudy (anonymous):

CAN U PLEASE TELL ME HOW U GOT THIS ANSWER

OpenStudy (karatechopper):

he plugged the numbers in

OpenStudy (karatechopper):

do u kno (x,y) is the way a coordinate is written?

OpenStudy (karatechopper):

am i right saif?

OpenStudy (anonymous):

YES

OpenStudy (karatechopper):

did i ask u?! jk:)

OpenStudy (saifoo.khan):

y - 3 = 3/5 x - 2

OpenStudy (saifoo.khan):

Sorry i was in another tab.

OpenStudy (saifoo.khan):

y - 3 = 3/5 x - 2

OpenStudy (saifoo.khan):

Yes, i did just inserted the asnwers.

OpenStudy (saifoo.khan):

y - 3 = 3/5 x - 2

OpenStudy (saifoo.khan):

i mean numbers*

OpenStudy (karatechopper):

im right! and saif can u look at the post i made out to u in efc?

OpenStudy (karatechopper):

i need one medal until im in level 40!!!

OpenStudy (anonymous):

THANKS

OpenStudy (karatechopper):

o thnx

OpenStudy (saifoo.khan):

kfc, u kicked me out from there.